OpenStudy (anonymous):

Which of the following properties tells us that 3x+(2y+4)=(3x+2y)+4 ? A. Commutative Property of Addition B. Associative Property of Addition C. Distributive Property D. None of the above

OpenStudy (zehanz):

Dear @brittanyburton, you and I both know you asked the same kind of question only 10 minutes ago. I did my best to give you an explanation. I really am convinced that if you would read what I explained, you would be able to answer this one yourself.

OpenStudy (anonymous):

LOL no i need the answer

OpenStudy (anonymous):

I put A but I need to know if its right okay

OpenStudy (zehanz):

a+b=b+a is commutative, a+(b+c)=(a+b)+c is associative, so it's B I'm afraid...